PET has a Young's, Modulus of 3.5 GPa and PLGA 1.2 GPa. A polymer blend is prepared with 45:55 mixture (PET:PlGA) . What is the Young's, Modulus of this blend (use mixture rule). Calculate the percent load borne by PET when the blend is subjected to a load. If after 10 days, entire PET had degraded, what is the Young's Modulus of this blend.

1
Expert's answer
2022-04-14T15:04:04-0400

Ym = YPET × 0.45 + YPIGA × 0.55 = (3.5 GPa) × 0.45 + (1.2 GPa) × 0.55 = 2.24 GPa

Total elongation = elongation of 1 + elongation of 2

so here Youngs modulues for mixture after 10 days as per rule of degradation

Ym = 3.5/2 + 1.2/3 = 2.14 GPa